OpenStudy (anonymous):

As the limit "phata" approaches pie over 6. of (phata)^2 tan(phata). Find f(x).

OpenStudy (aum):

I don't know what "phata" is. I have heard of phi and theta but not phata. But the problem does not change if I use theta instead. \[\large \lim_{\theta \rightarrow \frac{\pi}{6}}(\theta)^2\tan(\theta) = \left( \frac{\pi}{6} \right)^2\tan \left(\frac{\pi}{6}\right) = ? \]

OpenStudy (anonymous):

graphic ...